# Vector-Quantized Vision Foundation Models for Object-Centric Learning

### v1

@@ 3.2 Fully Utilize VFMs in OCL
37.9 $\pm$2.1 81.7 $\pm$2.6 29.3 $\pm$1.1 47.8 $\pm$1.4 37.2 $\pm$0.7 38.3 $\pm$0.9 7.3 $\pm$0.1 27.8 $\pm$0.9 17.8 $\pm$0.1 24.1 $\pm$0.3 14.1 $\pm$0.1 36.1 $\pm$0.3
Table 1: Object discovery performance with DINO2 ViT (s/14) for OCL encoding. VVO is instantiated as VQDINO; Tfd, Tfdt, MLP and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
resolution= ClevrTex #slot=11 COCO (panoptic) #slot=7 VOC #slot=6 MOVi-D #slot=21, conditional
256$\times$256 ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO ARI ARIfg mIoU mBO
SLATE-SAM 12.4
$\pm$1.1
80.3
$\pm$0.6
28.4
$\pm$0.3
44.0
$\pm$0.8
34.7
$\pm$0.6
35.8
$\pm$1.4
6.4
$\pm$0.3
25.9
$\pm$0.4
12.7
$\pm$0.4
19.5
$\pm$0.9
11.7
$\pm$0.2
31.6
$\pm$1.1
@@
16.7 $\pm$0.3 76.9 $\pm$0.3 32.5 $\pm$0.7 49.2 $\pm$0.9 35.7 $\pm$0.6 36.8 $\pm$0.6 7.2 $\pm$0.0 27.1 $\pm$0.4 16.9 $\pm$1.3 19.6 $\pm$1.4 13.6 $\pm$0.2 36.0 $\pm$0.9
Table 2: Object discovery performance with SAM2 Hiera+FPN (t/16) for OCL decoding. VVO is instantiated as VQSAM; Tfd, Tfdt, MLP and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
4 Experiment
@@
4.1 Set up the Benchmark
Datasets. We include both synthetic and real-world datasets. ClevrTex [^2] https://www.robots.ox.ac.uk/~vgg/data/clevrtex comprises synthetic images, each with about 10 geometric objects scattered in complex background. MOVi-D [^3] https://github.com/google-research/kubric/blob/main/challeng
es/movi/README.md#movi-d
contains synthetic videos, each with up to 20 daily objects dropping and bumping. COCO [^4] https://cocodataset.org is a recognized real-world image dataset, and we use its challenging panoptic segmentation. VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC is a real-world image dataset, and we use its instance segmentation. We also report results on real-world video dataset YTVIS [^6] https://youtube-vos.org/dataset/vis version HQ [^7] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube. We choose Physion [^8] https://physion-benchmark.github.io for visual prediction and reasoning as it contains common object interactions, requiring algorithms to learn dynamics like support, roll and link, then to predict and reason about future scene states.
@@ 4.1 Set up the Benchmark
Comparison. Instead of copying existing results, we reproduce all baselines to realize fair comparison. As detailed in Appendices A and B, we use identical data augmentation, VFMs in OCL encoding and training recipes for all experiment items unless not applied. We instantiate all baselines’ VAE part as TAESD [^9] https://huggingface.co/docs/diffusers/en/api/models/autoenco
der_tiny
, which is a large-scale pretrained StableDiffusion [^10] https://huggingface.co/spaces/stabilityai/stable-diffusion module, to build all strong baselines.
ARI ARIfg mIoU mBO
Using higher resolution: 384$\times$384
resolution=384$\times$384 COCO (panoptic) #slot=7
SLATE-DINO 50.9
$\pm$0.3
51.0
$\pm$0.9
9.0
$\pm$0.1
27.9
$\pm$0.7
@@
resolution=256$\times$256 COCO (panoptic) #slot=7
SLATE-DINO-SlotAttention 34.1
$\pm$0.7
34.9
$\pm$0.6
7.0
$\pm$0.1
26.2
$\pm$0.9
VQDINO Tfd -SlotAttention 37.7 $\pm$0.5 39.3 $\pm$0.4 7.7 $\pm$0.1 28.0 $\pm$0.6
SLATE-DINO-BO-QSA 34.4
$\pm$0.3
35.4
$\pm$0.5
7.1
$\pm$0.1
@@
$\pm$0.7
VQDINO Tfd -BO-QSA 38.5 $\pm$0.4 39.6 $\pm$0.4 7.8 $\pm$0.0 28.5 $\pm$0.4
Table 3: VVO using higher resolution (upper) and different aggregators (lower) on object discovery. By default, we use BO-QSA for all our experiment items, including the baselines.
ARI ARIfg mIoU mBO
@@
resolution=256$\times$256 COCO (panoptic) #slot=7
@@
resolution=256$\times$256 COCO (panoptic) #slot=7
SPOT-DINO 37.5
$\pm$0.4
38.7
$\pm$0.3
7.8
$\pm$0.1
28.1
$\pm$0.9
@@
+self-training
39.3 $\pm$0.3 40.4 $\pm$0.4 8.0 $\pm$0.0 29.6 $\pm$0.5
resolution=256$\times$256 YTVIS (HQ) #slot=7, unconditional
VideoSAUR-DINO 34.0
$\pm$0.6
39.1
$\pm$0.9
27.2
$\pm$0.4
32.2
$\pm$0.6
@@
Smd
40.7 $\pm$0.5 43.5 $\pm$0.6 29.2 $\pm$0.2 35.1 $\pm$0.5
Table 4: VVO compared with general improvers (upper) and SotA methods (lower) on object discovery. SPOT uses both Transformer decoder with 9 permutations Tfd9 and “self-training”; VideoSAUR uses SlotMixer Smd as its decoder.
4.2 Evaluate on Object Discovery
Object discovery task intuitively shows how well those slots separate different objects. We evaluate the byproduct object segmentation accuracy with Adjusted Rand Index [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metri
cs.adjusted_rand_score.html
(ARI), ARIfg (foreground), mean Intersection-over-Union [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metri
cs.jaccard_score.html
(mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO).
@@ 4.3 Evaluate on Set Prediction
0.64 $\pm$0.1 0.59 $\pm$0.1
Table 5: Set prediction performance on COCO (panoptic, #slot=7).
Set prediction task directly shows how much object information those slots grasp. We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023). We measure top1 accuracy of the classified class labels while we measure the R2 score [^13] https://scikit-learn.org/stable/modules/generated/sklearn.metri
cs.r2_score.html
of the regressed bounding box coordinates.
As shown in Table 5, compared with baseline DINOSAUR, our VVO, i.e., VQDINO-MLP, obtains better object classification and bounding box regression. Thus, our models’ object representations are of higher quality.
4.4 Deploy to the Downstream
Figure 5: Visual prediction (upper) and reasoning (lower) performance on Physion (#slot=8). VVO has smaller prediction error in all time steps, and higher reasoning accuracy in later time steps.
Better object representation benefits downstream tasks. We follow the convention to pretrain OCL models on Physion and represent this dataset as slots. Then the object-centric dynamics model SlotFormer (Wu et al., 2023a) is trained on those slots in an auto-regressive manner along the time dimension. We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO-MLP to extract slots.
@@ 4.5 Ablate the Architecture
shared DINO2 ViT DINO2 ViT 38.5
$\pm$0.4
39.6
$\pm$0.4
separate TAESD encoder ResNet18 28.3 $\pm$1.3 29.1 $\pm$1.5
TAESD encoder DINO2 ViT 34.6
$\pm$0.7
35.8
$\pm$0.5
SAM2 encoder DINO2 ViT 34.9
$\pm$0.3
35.8
$\pm$0.4
Table 6: VVO’s two key designs: (i) Using VFM representation for encoding is better than using non-VFMs; (ii) Sharing the OCL encoder as VAE encoder to obtain targets is better than using separate VAE and OCL encoders. Results are on COCO panoptic.
@@
Table 7: VVO’s VQ variant: All our three tricks are beneficial to the overall performance boosts. In comparison to VVO’s key designs, these tricks are more like the cherry on top. Results are on COCO panoptic with settings consistent with the above.
@@
Figure 6: Effects of tricks in our VQ variant: (upper) Gumbel noise improves codebook utilization, where “CV” means Coefficient of Variation, and curves are smoothed by Gaussian kernel of size 10; (lower) Annealing residual connection improves VAE pretraining, and the blue curve’s turning point at epoch 4 is where the residual connection anneals to zero. Results are from the VAE pretraining of VQDINO-Dfz on COCO panoptic.
Conclusion
@@
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Dataset Processing
The datasets for image OCL are ClevrTex, COCO (panoptic) and VOC. The datasets for video OCL are MOVi-D and YTVIS. For different datasets, we have mostly the same processing but a bit differences.
A.1 Shared Preprocessing
To accelerate experiments, we adopt the dataset conversion scheme: Convert all datasets into LMDB database format, and load it into NVMe disk or RAM disk to minimize I/O overhead and maximize throughput.
Specifically, all images or video frames are resized by short side to 256, as well as their segmentation masks, and then are stored data type uint8 to save space. For images or videos, we use default interpolation mode; but for segmentation masks, we use NEAREST-EXACT interpolation mode.
During training, for both images and videos, we use spatial resolution of 256$\times$256 as inputs and outputs. We normalize the input images or videos by ImageNet-Mean and -Std, which are the common normalization scheme using by VFMs like DINO and SAM. For videos, we also using random strided temporal cropping of fixed window size 6 to accelerate training and improve generalization.
During testing, there is no difference to that of training for images. But for videos, we need to remove the random strided temporal cropping and use the full video of 24 time steps.
A.2 Dataset-Specific Processing
ClevrTex. As there are 10 objects at most in each image, we use number of slots queries of 10 + 1, where 1 represents for the background.
COCO & VOC. We use the panoptic segmentation annotations of COCO for evaluation, which is more challenging than instance segmentation. Besides, OCL’s queries-based workflow intuitively fits panoptic segmentation rather than instance segmentation or semantic segmentation. There can be lots of or only two objects in an image, but we just follow the popular hyperparameter selection, i.e., 7 slots for COCO and 6 slots for VOC, respectively.
MOVi-D. There are abundant annotations like depth, optical-flow, camera intrinsic/extrinsic parameters and so on. But for simple comparison, we only utilize the segmentation masks for evaluation and the bounding boxes as slots initialization. As there are 23 objects in these dataset, we use number of slots queries of 23+1.
YTVIS. We use the original YTVIS videos and HQ-YTVIS segmentation annotations. The latter provides high-quality masks. During VAE pretraining, we utilize all the video frames. But during OCL training, we only utilize the every-5-time-step annotations for calculating the performance.
Appendix B Training Recipe
Following the convention, there are two training stages. The pretraining stage is for training VAE modules on corresponding datasets, so as to learn good discrete intermediate representation. The OCL training stage, which utilize the pretraining VAE representation as guidance for object centric learning.
B.1 Pretrain VAE
For all datasets, we use total number of training iterations 30,000, and validation interval iterations 600, so that we have about 50 checkpoints for every OCL model on every dataset. But to save storage, we save only the latter half 25 checkpoints. This is unified across all datasets.
The batch size for image datasets is 64, while for video datasets it is 16. This holds for both training and validation. This setting is shared across all datasets.
For multi-process, we use the number of workers 4. This holds for both training and validation. This setting is shared across all datasets.
We use Adam optimizer with initial learning rate of 2e-3. And the learning rate is manipulated by cosine annealing scheduling, with a linear warmup during first 1/20 of total steps. This setting is shared across all datasets.
We use automatic mixed precision provided by PyTorch autocast API plus compile. Along with this, we use PyTorch builtin gradient scaler, so as to apply gradient clipping of maximum norm 1.0. This setting is shared across all datasets.
B.2 Train OCL Model
For all datasets, we use total number of training iterations 5,$0$, and validation interval iterations 1000, so that we have about 50 checkpoints for every OCL model on every dataset. But to save storage, we only save the latter half of the total checkpoints. This setting is shared across all datasets.
The batch size for image datasets is 32 for both training and validation, while for video datasets it is 8 for training and 4 for validation as there are more time steps in videos during validation. This setting is shared across all datasets.
For multi-process, we use the number of workers 4. This holds for both training and validation. This setting is shared across all datasets.
The objective is the CE and MSE loss following the basis methods original design.
Metrics used here are ARI, mIoU and mBO, calculating the panoptic segmentation accuracy of both objects and the background. This setting is shared across all datasets.
We use Adam optimizer with initial learning rate of 2e-4. And the learning rate is manipulated by cosine annealing scheduling, with a linear warmup during first 1/20 of total steps. This setting is shared across all datasets.
We use auto mixed precision provided by PyTorch autocast API and compile. Along with this, we use PyTorch builtin gradient scaler, so as to apply gradient clipping of maximum norm 1.0 for images and 0.02 for videos.

### v3

@@ 3.3. Compare Architectures
72.2 ±0.2 81.9
±2.0
57.6 ±0.7 56.8 ±0.8 18.3 ±0.4 28.7
±1.0
27.2 ±0.1 25.8 ±0.1 19.1 ±0.5 24.2 ±0.3 36.9 ±0.1 35.9 ±0.2
Table 1. Object discovery performance with DINO2 ViT (s/14) for OCL encoding. VVO is instantiated as VQDINO; Tfd, TfdT, Mlp and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
resolution= ClevrTex #slot=11 COCO #slots=7 VOC #slot=6 MOVi-D #slot=21 conditional
256$\times$256 ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU ARI ARIfg mBO mIoU
SLATE-SAM 18.0
±0.3
82.8
±0.9
51.8
±0.8
50.6
±0.8
12.5
±0.8
27.4
±1.4
22.7
±0.8
21.5
±0.8
16.3
±0.7
21.4
±1.3
34.8
±0.5
33.9
±0.6
@@
30.0 ±0.3 78.6
±0.8
57.3 ±0.2 56.1 ±0.2 17.3 ±0.4 26.0
±0.2
27.2 ±0.4 25.8 ±0.4 17.3 ±0.5 20.2 ±2.3 36.3 ±0.8 35.1 ±0.9
Table 2. Object discovery performance with SAM2 Hiera+FPN (t/16) for OCL decoding. VVO is instantiated as VQSAM; Tfd, TfdT, Mlp and Dfz are Transformer, Transformer-temporal, MLP and Diffusion for OCL decoding respectively.
4. Experiment
@@
4.1. Set up the Benchmark
Datasets. We include both synthetic and real-world datasets. ClevrTex [^2] https://www.robots.ox.ac.uk/~vgg/data/clevrtex comprises synthetic images, each with about 10 geometric objects scattered in complex background. MOVi-D [^3] https://github.com/google-research/kubric/blob/main/challenges/movi/README.md#
movi-d
contains synthetic videos, each with up to 20 daily objects dropping and bumping. COCO [^4] https://cocodataset.org is a recognized real-world image dataset, and we use its challenging panoptic segmentation. VOC [^5] http://host.robots.ox.ac.uk/pascal/VOC is a real-world image dataset, and we use its instance segmentation. We also report results on real-world video dataset YTVIS version HQ [^6] https://github.com/SysCV/vmt?tab=readme-ov-file#hq-ytvis-high-quality-video-instance-segmentation-dataset, which contains large-scale short videos from YouTube. We choose Physion [^7] https://physion-benchmark.github.io for visual prediction and reasoning as it contains common object interactions, requiring algorithms to learn dynamics like support, roll and link, then to predict and reason about future scene states.
@@ 4.1. Set up the Benchmark
Comparison. Instead of copying existing results, we reproduce all baselines to realize fair comparison. We use identical data augmentation, VFMs in OCL encoding and training recipes for all experiment items unless not applied. We instantiate all baselines’ VAE part as TAESD [^8] https://huggingface.co/docs/diffusers/en/api/models/autoencoder_tiny, which is a large-scale pretrained StableDiffusion [^9] https://huggingface.co/spaces/stabilityai/stable-diffusion module, to build all strong baselines.
ARI ARIfg mBO mIoU
Using higher resolution: 384$\times$384 (336)
resolution=384$\times$384 (336) COCO #slot=7
SLATE-DINO 41.4
±1.0
34.0
±0.3
27.4
±0.4
25.9
±0.5
@@
resolution=256$\times$256 (224) COCO #slot=7
SLATE-DINO-SlotAttention 17.0
±1.3
28.3
±0.5
26.4
±0.4
25.1
±0.3
VQDINO Tfd -SlotAttention 20.8 ±2.0 31.5 ±1.2 29.4 ±0.9 27.9 ±1.1
SLATE-DINO-BO-QSA 17.5
±0.6
28.8
±0.3
@@
±0.3
25.4
±0.3
VQDINO Tfd -BO-QSA 21.1 ±2.1 31.5 ±1.1 29.6 ±0.7 28.2 ±0.8
Table 3. VVO using higher resolution (upper) and different aggregators (lower) on object discovery. By default, we use BO-QSA for all our experiment items, including the baselines.
ARI ARIfg mBO mIoU
@@
resolution=256$\times$256 (224) COCO #slot=7
@@
resolution=256$\times$256 (224) COCO #slot=7
SPOT-DINO 20.3
±0.7
41.1
±0.3
30.4
±0.1
29.0
±0.9
@@
21.3 ±0.4 42.3 ±1.0 31.4 ±0.2 29.9 ±0.3
resolution=256$\times$256 (224) YTVIS (HQ) #slot=7, unconditional
VideoSAUR-DINO 33.0
±0.6
49.0
±0.9
30.8
±0.4
30.1
±0.6
@@
SmdT
35.7 ±0.5 49.5 ±0.6 32.7 ±0.2 31.6 ±0.5
Table 4. VVO versus general improvers (upper) and SotA methods (lower) on object discovery. SPOT uses Transformer with 9 permutations Tfd9 as decoder while VideoSAUR uses SlotMixer SmdT.
4.2. Evaluate on Object Discovery
Object discovery task intuitively shows how well those slots separate different objects. We evaluate the byproduct object segmentation accuracy with Adjusted Rand Index [^10] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.adjusted_rand_sco
re.html
(ARI), ARIfg (foreground), mean Intersection-over-Union [^11] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.jaccard_score.html (mIoU) and mean Best Overlap (Uijlings et al., 2013) (mBO).
@@ 4.3. Evaluate on Set Prediction
0.64 ±0.1 0.59 ±0.1
Table 5. Set prediction performance on COCO (panoptic, #slot=7).
Set prediction task directly shows how much object information those slots grasp. We use OCL to represent dataset COCO as slots, and use a small MLP to predict the object class label and bounding box corresponding to each slot by following this work (Seitzer et al., 2023). We measure top1 accuracy of the classified class labels while we measure the R2 score [^12] https://scikit-learn.org/stable/modules/generated/sklearn.metrics.r2_score.html of the regressed bounding box coordinates.
As shown in Table 5, compared with baseline DINOSAUR, our VVO, i.e., VQDINO Mlp , obtains better object classification and bounding box regression. Thus, our models’ object representations are of higher quality.
4.4. Deploy to the Downstream
Figure 5. Visual prediction (upper) and reasoning (lower) performance on Physion (#slot=8). VVO has smaller prediction error in all time steps, and higher reasoning accuracy in later time steps.
Better object representation benefits downstream tasks. We follow the convention to pretrain OCL models on Physion and represent this dataset as slots. Then the object-centric dynamics model SlotFormer (Wu et al., 2023a) is trained on those slots in an auto-regressive manner along the time dimension. We use temporal versions of DINOSAUR and our VVO, i.e., VQDINO Mlp to extract slots.
@@ 4.5. Ablate the Architecture
shared DINO2 ViT DINO2 ViT 21.1
±2.1
31.5
±1.1
separate TAESD encoder ResNet18 15.4 ±3.3 24.1 ±2.5
TAESD encoder DINO2 ViT 17.5
±0.6
28.8
±0.3
SAM2 encoder DINO2 ViT 18.3
±1.3
29.1
±0.9
Table 6. VVO’s two key designs: (i) Using VFM representation for encoding is better than using non-VFMs; (ii) Sharing the OCL encoder as VAE encoder to obtain targets is better than using separate VAE and OCL encoders. Results are on COCO panoptic.
@@
Table 7. VVO’s VQ variant: All our three tricks are beneficial to the overall performance boosts. In comparison to VVO’s key designs, these tricks are more like the cherry on top. Results are on COCO panoptic with settings consistent with the above.
@@
Figure 6. Effects of tricks in our VQ variant: (upper) Gumbel noise improves codebook utilization, where “CV” means Coefficient of Variation, and curves are smoothed by Gaussian kernel of size 10; (lower) Annealing residual connection improves VAE pretraining, and the blue curve’s turning point at epoch 4 is where the residual connection anneals to zero. Results are from the VAE pretraining of VQDINO
Dfz
on COCO panoptic.
5. Analysis
We provide both mathematical modeling and statistical verification on why and how our VVO is better.
5.1. Aggregation as Clustering
Figure 7. Better objectness helps OCL aggregation. Green and orange areas stand for objects $\bm{o}_{1}$ and $\bm{o}_{2}$, where $\bm{v}_{1}$ and $\bm{v}_{2}$ are arbitrary super-pixels and $\bm{s}_{*}$ is $\bm{o}_{2}$’s centroid. But the actual query $\bm{s}\sim\mathcal{N}(\bm{s}_{*},\bm{\sigma}^{2})$. In VFM super-pixel space, the distance $d_{12}$ between $\bm{v}_{1}$ and $\bm{v}_{2}$ is larger, i.e., better objectness, thus $\bm{s}$ has higher probability $p(d_{s1}>d_{s2})$ to represent $\bm{o}_{2}$ correctly, compared with that in non-VFM super-pixel space.
Mathematic Modeling
(14) $d_{*1}=d(\bm{s}_{*},\bm{v}_{1})>d(\bm{s}_{*},\bm{v}_{2})=d_{*2}$
[…]
According to Figure 1 the first observation, in VFM super-pixel space, points of the same object have smaller distances while points from different objects have larger distances, compared to that in non-VFM space. This means the separation plane is closer to $\bm{s}_{*}$ in non-VFM space. Therefore, $p_{2}$ in VFM space is bigger than in non-VFM space. $\square$
Statistical Verification
Equation 15 is very difficult to solve, but we can use Monte Carlo estimation:
(16) $p_{2}=\frac{1}{N}\Sigma_{i=1}^{N}\mathbbm{1}[\bm{v}_{i}\in\bm{o}_{2}]$
Specifically, we respectively use ResNet18, DINO2 ViT and SAM2 encoder to extract images into super-pixels, then apply KMeans clustering to find each clusters mean and std. From the Gaussian distributions defined by these mean and std, we repeatedly draw samples and calculate the ratio that such samples fall within their original clusters. This ratio is the correct aggregation probability.
Figure 8. Statistical correlation between correct aggregation probability and object discovery accuracy on VOC.
A step further, we can calculate the statistical correlation between the correct aggregation probability and the corresponding OCL performance. As shown in Figure 8, the correlation is roughly significant. Considering that correct aggregation probability can be calculated before the expensive OCL training, our Equation 15 is instructive to training better OCL models by selecting OCL encoders.
5.2. Shared Quantization and Optimization Noise
Mathematic Modeling
(17) $\frac{\partial\mathrm{MSE}(\bm{Q}^{\prime},\mathrm{sg}(\bm{Q}))}{\partial\bm{Z}}=2(\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}))\frac{\partial\bm{Q}^{\prime}}{\partial\bm{Z}}$
(18) $\mathbb{E}[\bm{Q}^{\prime}]=\mathbb{E}[f(\bm{Z})]\approx\bm{Q}$
(19) $\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q})]\approx 0$
But if instead of sharing $\bm{\phi}_{\mathrm{e}}$, we use an extra VAE encoder plus $\bm{\phi}_{\mathrm{q}}$ to obtain the target, denoted as $\bm{Q}_{2}$, then $\bm{Q}_{2}\neq\bm{Q}$ according to Figure 1 the second observation. Substitute $\bm{Q}_{2}$ into Equation 17 and the residual error $\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})$ would be systematically biased:
(20) $\mathbb{E}[\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}_{2})]\neq 0$
which increases noise in optimization.
Under CE loss, the gradient with respective to $\bm{Z}$ is:
(21) $\frac{\partial\mathrm{CE}(\bm{Q}^{\prime},\mathrm{sg}(\bm{Q}))}{\partial\bm{Z}}=\frac{\partial f(\bm{Z})^{T}}{\partial\bm{Z}}(\bm{Q}^{\prime}-\mathrm{sg}(\bm{Q}))$
where $\frac{\partial f(\bm{Z})}{\partial\bm{Z}}$ is the Jacobian matrix of $f(\bm{Z})$ with respect to $\bm{Z}$. Anyway, this has similar structure to Equation 17 and thus does not alter our judgment above. $\square$
Statistical Verification
Figure 9. Gradient distributions tend to disperse or have outliers if not using our shared VFM feature quantization. Results here are collected at time step 10k, on dataset VOC.
The statistical bias in Equation 20 manifests as the gradient distribution of different layers, especially that of slots. No statistical bias is usually associated with a smooth zero-mean small-std Gaussian gradient distribution.
Our technique, shared VFM feature quantization as reconstruction targets, includes two key points, sharing and quantization. Thus we compare three cases: (i) Shared and quantized, where DINO features are both used for aggregation and quantized for reconstruction; (ii) Shared but not quantized, where DINO features are used for aggregation and direct for reconstruction without quantization; (iii) Not shared but quantized, where DINO features are used for aggregation while quantized SAM features are used for reconstruction.
As shown in Figure 9, using CE or MSE as the loss, our shared quantization achieves smoother Gaussian gradient distributions, which benefits OCL training. If not shared but quantized, then gradients tend to disperse; while if shared but not quantized, then gradients have some outliers.
6. Conclusion